Math 312 Spring 98

Quiz 2

Name:

Complex Analysis

Date:  Friday January 23, 1998
Time Begun:  Ron Buckmire

Time Ended:

Topic : Solutions of Equations of a Complex Numbers

The point of this quiz is to test your ability to solve equations which involve complex variables, and to understand the multiple forms of complex numbers.

Instructions:

1. Once you open the quiz, you have 60 minutes to complete it.

2. You may use the book or any of your class notes. You must work alone.

3. If you use your own paper, please staple it to the quiz before coming to class. If you don’t have a stapler, buy one.

4. After completing the quiz, sign the pledge below stating on your honor that you have adhered to these rules.

5. Your solutions must have enough details such that an impartial observer can read your work and determine HOW you came up with your solution.

6. Relax and enjoy...

7. This quiz is due on Monday, January 26, in class. NO LATE QUIZZES WILL BE ACCEPTED.

Pledge: I, , pledge my honor as a human being and Occidental student, that I have followed all the rules above to the letter and in spirit.

Math 312 Spring 98  Quiz 2

SHOW ALL YOUR WORK

1. Find all the solutions of z^{2} + (2i − 3)z + 5 − i = 0 and write them in rectangular form.

2. Show that z_{1} = 2 − 3i and z_{2} = 1 + i solve the above equation and write them in exponential form.